# Eliminating Biased Length Reliance of Direct Preference Optimization via Down-Sampled KL Divergence

### v1

@@ Abstract
Direct Preference Optimization (DPO) has emerged as a prominent algorithm for the direct and robust alignment of Large Language Models (LLMs) with human preferences, offering a more straightforward alternative to the complex Reinforcement Learning from Human Feedback (RLHF). Despite its promising efficacy, DPO faces a notable drawback: “verbosity”, a common over-optimization phenomenon also observed in RLHF. While previous studies mainly attributed verbosity to biased labels within the data, we propose that the issue also stems from an inherent algorithmic length reliance in DPO. Specifically, we suggest that the discrepancy between sequence-level Kullback–Leibler (KL) divergences between chosen and rejected sequences, used in DPO, results in overestimated or underestimated rewards due to varying token lengths. Empirically, we utilize datasets with different label lengths to demonstrate the presence of biased rewards. We then introduce an effective downsampling approach, named SamPO, to eliminate potential length reliance. Our experimental evaluations, conducted across three LLMs of varying scales and a diverse array of conditional and open-ended benchmarks, highlight the efficacy of SamPO in mitigating verbosity, achieving improvements of 5% to 12% over DPO through debaised rewards [^1] Our codes: https://github.com/LuJunru/SamPO/..
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a crucial strategy for effectively align Large Language Models (LLMs) with human minds Zhao et al. (2023a); Yang et al. (2023); Pan et al. (2023b), showcasing significant improvements of LLM’s instruct-following capability compared with the other two popular approaches: pre-training and supervised fine-tuning (SFT). In fact, a series of leading LLMs have adopted RLHF as the final stage of their entire training pipelines Ouyang et al. (2022); Achiam et al. (2023); Touvron et al. (2023); Bi et al. (2024); Bai et al. (2023).
@@ 3.1 Preliminary Background of DPO
$\nabla_{\theta}\mathcal{L}_{dpo}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}[\Delta\cdot\mathcal{M}]$ (5)
@@ 5 Experimental Results
In this section, following above designs, we first report the group experiments of length reliance (§ 5.1), then present comparison studies against strong baselines (§ 5.2). We discuss quantitative results in the main body. We leave more ablation studies and case analysis in Appendix G,  H, and  I.
@@ A.2 Gradients of Token-level DPO reward
$\nabla_{\theta}\mathcal{L}_{dpo}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}[\Delta\cdot\mathcal{M}]$
@@ Appendix B Evaluation Details
For all conditional benchmarks, we use a stable and popular evaluation framework “lm-evaluation-harness” Gao et al. (2023b) [^3] Official tool page of lm-eval: https://github.com/EleutherAI/lm-evaluation-harness. As for open-ended benchmarks, we report specific evaluation templates for AlpacaEval2, HH-RLHF and TL;DR in Appendix J. Particularly, we use the official tool to evaluate AlpacaEval2444https://github.com/tatsu-lab/alpaca_eval/. The version of GPT-4 evaluator is all set as: gpt-4-turbo.
@@ Appendix G Influence of Different Random Seed
Appendix I Case Analysis
In this section, we introduce exemplars to analyze the qualitative outputs from different approaches. We analyze two cases for each test set: AlpacaEval2 (§ I.1), HH-RLHF (§ I.2), and TL;DR (§ I.3), in each section respectively.
I.1 Case analysis on AlpacaEval2
@@
I.2 Case analysis on HH-RLHF
@@ I.3 Case analysis on TL;DR
Appendix J GPT-4 Judged Winner Templates

### v2

@@ Abstract
Direct Preference Optimization (DPO) has emerged as a prominent algorithm for the direct and robust alignment of Large Language Models (LLMs) with human preferences, offering a more straightforward alternative to the complex Reinforcement Learning from Human Feedback (RLHF). Despite its promising efficacy, DPO faces a notable drawback: “verbosity”, a common over-optimization phenomenon also observed in RLHF. While previous studies mainly attributed verbosity to biased labels within the data, we propose that the issue also stems from an inherent algorithmic length reliance in DPO. Specifically, we suggest that the discrepancy between sequence-level Kullback–Leibler (KL) divergences between chosen and rejected sequences, used in DPO, results in overestimated or underestimated rewards due to varying token lengths. Empirically, we utilize datasets with different label lengths to demonstrate the presence of biased rewards. We then introduce an effective downsampling approach, named SamPO, to eliminate potential length reliance. Our experimental evaluations, conducted across three LLMs of varying scales and a diverse array of conditional and open-ended benchmarks, highlight the efficacy of SamPO in mitigating verbosity, achieving improvements of 5% to 12% over DPO through debaised rewards [^1] Our code can be accessed at: https://github.com/LuJunru/SamPO/..
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a crucial strategy for effectively align Large Language Models (LLMs) with human minds Zhao et al. (2023a); Yang et al. (2023); Pan et al. (2023b), showcasing significant improvements of LLM’s instruct-following capability compared with the other two popular approaches: pre-training and supervised fine-tuning (SFT). In fact, a series of leading LLMs have adopted RLHF as the final stage of their entire training pipelines Ouyang et al. (2022); Achiam et al. (2023); Bi et al. (2024).
@@ 3.1 Preliminary Background of DPO
$\nabla_{\theta}\mathcal{L}_{dpo}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}[\beta\sigma(-\Delta)\mathcal{M}]$ (5)
@@ 5 Experimental Results
In this section, following above designs, we first report the group experiments of length reliance (§ 5.1), then present comparison studies against strong baselines (§ 5.2). We discuss quantitative results in the main body. We leave more ablation studies and case analysis in Appendix G,  H, and  J.
@@ 6 Conclusion
Acknowledgment
We thank Shiyue Xu for pointing out the error in Equation 5 in the previous draft.
@@ A.2 Gradients of Token-level DPO reward
$\nabla_{\theta}\mathcal{L}_{dpo}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}[\beta\sigma(-\Delta)\mathcal{M}]$
@@ Appendix B Evaluation Details
For all conditional benchmarks, we use a stable and popular evaluation framework “lm-evaluation-harness” Gao et al. (2023b) [^3] Official tool page of lm-eval: https://github.com/EleutherAI/lm-evaluation-harness. As for open-ended benchmarks, we report specific evaluation templates for AlpacaEval2, HH-RLHF and TL;DR in Appendix K. Particularly, we use the official tool to evaluate AlpacaEval2444https://github.com/tatsu-lab/alpaca_eval/. The version of GPT-4 evaluator is all set as: gpt-4-turbo.
@@ Appendix G Influence of Different Random Seed
MRC Logical Reasoning RolePlay Avg.
SFT Base 81.25 69.52 59.12 69.96
w/ DPO 85.33 73.25 57.41 72.00
w/ SamPO 87.50 83.57 63.61 78.23
Table 8: Human Evaluation results of a Qwen1.5-72B-based SFT model and its two further fine-tuned versions, applying with DPO and SamPO respectively.
@@
Appendix I Human Evaluation of SamPO
In addition to the aforementioned automated evaluation, we further conduct large-scale human evaluation to study the effectiveness of the SamPO algorithm when applied to super large LLM (e.g, over 50B). We use an LLM fine-tuned based on Qwen1.5-72B Bai et al. (2023) as a starting point and fine-tune it for 1 epoch using the proposed SamPO method. The training data is a general preference dataset of around 480,000 samples.
We report the results of human evaluation in Table 8, covering three most popular scenarios: general Machine Reading Comprehension (MRC), logical reasoning (e.g., math or logic questions), and open domain dialogues in role-play settings. We have hired a 30-person annotation team, each of whom has at least a bachelor’s degree or above. And each test scenario contains 500 to 1,000 carefully crafted challenging instances, which then cross-labeled by multiple professional annotators.
Our scoring criteria are relatively simple, distinguishing only between incorrect and acceptable responses, rather than employing a more fine-grained scoring system (such as a 1 to 10 point scale). We observe that SamPO significantly outperforms both the SFT Base and DPO methods on all three tasks.
Appendix J Case Analysis
In this section, we introduce exemplars to analyze the qualitative outputs from different approaches. We analyze two cases for each test set: AlpacaEval2 (§ J.1), HH-RLHF (§ J.2), and TL;DR (§ J.3), in each section respectively.
J.1 Case analysis on AlpacaEval2
@@ Appendix I Human Evaluation of SamPO
J.2 Case analysis on HH-RLHF
@@
Appendix K GPT-4 Judged Winner Templates